OpenStudy (anonymous):

@dumbcow A recipe that makes 3 servings calls for 3/5 cup of apple juice and 3/4 cup of cranberry juice. How many cups of juice are there per serving? A. 4 and1/20 B. 1 and 7/20 C.2/3 D.9/20

OpenStudy (dumbcow):

add 3/5 + 3/4 then divide by 3

OpenStudy (dumbcow):

3/5 = 12/20 3/4 = 15/20